OpenStudy (anonymous):

Suppose that $2300 is initially invested in an account at a fixed interest rate, compounded continuously. Suppose also that, after two years, the amount of money in the account is $2517 . Find the interest rate per year

OpenStudy (anonymous):

your job is to solve \[2517=2300e^{2r}\] for r

OpenStudy (anonymous):

im not sure how to get the r..i have that exact formula written down

OpenStudy (anonymous):

from there im lost

OpenStudy (anonymous):

divide by 2300, take the log, divide by 2 \[\frac{2571}{2300}=e^{2r}\] \[\ln(\frac{2571}{2300})=2r\]

OpenStudy (anonymous):

\[r=\ln(\frac{2571}{2300})\div2\]

OpenStudy (anonymous):

or \[\frac{1}{2}\ln(\frac{2571}{2300})\]

OpenStudy (anonymous):

0.5 about right? it came to like .045 but it wants it rounded to the nearest hundreth

OpenStudy (anonymous):

and thank you so much btw!! you are a lifesaver

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

but don't round .045 to .5! leave at . 045 = 4.5%

OpenStudy (anonymous):

well thats thousandth place..so it would be .05

OpenStudy (anonymous):

Write your answer as a percentage. Do not round any intermediate computations, and round your percentage to the nearest hundredth.

OpenStudy (anonymous):

actually i get .05569

OpenStudy (anonymous):

oh wait i copied wrong

OpenStudy (anonymous):

lol so its .05 right?

OpenStudy (anonymous):

then yes =) thank you so much!!!!!!! =) i appreciate it!

OpenStudy (anonymous):

oh it says round percentage to nearest 100 so be careful again!

OpenStudy (anonymous):

yea it would be .05 =)

OpenStudy (anonymous):

ooh no

OpenStudy (anonymous):

??? what

OpenStudy (anonymous):

. 0450=4.5%

OpenStudy (anonymous):

it doesnt say round decimal. it says round percent

OpenStudy (anonymous):

ok..so 4.5 it is lol

OpenStudy (anonymous):

right? like at the bank 4.5% is a lot different than 5%

OpenStudy (anonymous):

yes lol...ok thank you =) ill b asking more questions in the coming of days

OpenStudy (anonymous):

k see you then